OpenStudy (waheguru):

How do i solve X^-4 = 1/16

OpenStudy (asnaseer):

do you know what \(x^{-4}\) means?

OpenStudy (waheguru):

Yea 1/x^4?

OpenStudy (asnaseer):

correct, so you have:\[ x^{-4}=\frac{1}{16}\\ \frac{1}{x^4}=\frac{1}{16}\]

OpenStudy (asnaseer):

can you solve now?

OpenStudy (waheguru):

So, what do i do from here?

OpenStudy (asnaseer):

come the two sides of the equation - what can you deduce from that?

OpenStudy (waheguru):

1*16=x^4*1?

OpenStudy (asnaseer):

correct

OpenStudy (asnaseer):

so you now know:\[x^4=16\]

OpenStudy (waheguru):

yea

OpenStudy (asnaseer):

can you write 16 as 2 to the power of something?

OpenStudy (waheguru):

I dont know...

OpenStudy (asnaseer):

what is the square root of 16?

OpenStudy (waheguru):

4?

OpenStudy (asnaseer):

good, so you know:\[16=4^2\]net, what is the square root of 4?

OpenStudy (asnaseer):

*next

OpenStudy (waheguru):

2

OpenStudy (waheguru):

what happened to the "x"

OpenStudy (asnaseer):

good, so \[4=2^2\]which means:\[16=4^2=(2^2)^2=2^4\]agreed so far?

OpenStudy (waheguru):

Yes, but I do not understand why we took out x

OpenStudy (asnaseer):

now we go back to see that:\[x^4=16=2^4\]

OpenStudy (asnaseer):

can you /see/ what the answer is now?

OpenStudy (waheguru):

2^4

OpenStudy (asnaseer):

if:\[a^4=b^4\]then we can say that a=b so, if:\[x^4=2^4\]then x=?

OpenStudy (waheguru):

ooh, x = 2

OpenStudy (waheguru):

U ROCK

OpenStudy (asnaseer):

:) there are actually more answers but I am not sure what level of maths you are at

OpenStudy (waheguru):

Grade 10

OpenStudy (asnaseer):

how old is that? I am in the UK so am not familiar with grade 10

OpenStudy (asnaseer):

for example, have you studied complex numbers yet?

OpenStudy (waheguru):

The answer you gave me was enough because we are learning quadratic relations

OpenStudy (agent0smith):

You probably want to at least include x= -2, since (-2)^4 = 16.

OpenStudy (asnaseer):

x=-2 is also an answer

OpenStudy (waheguru):

SO was the answer 2 or -2?

OpenStudy (asnaseer):

because:\[(-2)^4=16\]

OpenStudy (waheguru):

How would I properly show me work when solving this

OpenStudy (asnaseer):

when you got to the step:\[x^4=16\]you could proceed as follows:\[ x^4=16\\ x=\sqrt[4]{16}=\pm2\]

OpenStudy (waheguru):

Becase I have to show how I removed x

OpenStudy (asnaseer):

does that make sense?

OpenStudy (waheguru):

When I root a number to 4 in my calclulator I get 16?

OpenStudy (waheguru):

a number = 16

OpenStudy (asnaseer):

you are probably entering it incorrectly on your calc

OpenStudy (asnaseer):

enter 16 then press the square root key then press the square root key again this is because the fourth root of a number is the same as taking its square root twice

OpenStudy (waheguru):

Im getting 8 now?

OpenStudy (asnaseer):

hmmm.... all I can say is that you are not using the calc you have correctly are you aware of the wolframalpha web site?

OpenStudy (waheguru):

yea, I was just checking on that

OpenStudy (waheguru):

http://www.wolframalpha.com/input/?i=4%E2%88%9A16

OpenStudy (waheguru):

still 16...

OpenStudy (waheguru):

and √√ gets 8

OpenStudy (asnaseer):

what you entered there was 4 times the square root of 16

OpenStudy (waheguru):

Oh, so what would be the correct way

OpenStudy (asnaseer):

try this: http://www.wolframalpha.com/input/?i=sqrt%28sqrt%2816%29%29

OpenStudy (waheguru):

Thankyou very much, you have taught me so much today

OpenStudy (waheguru):

You are just an amazing teacher, thaks for your time

OpenStudy (asnaseer):

you can also do it like this: http://www.wolframalpha.com/input/?i=16^%281%2F4%29

OpenStudy (asnaseer):

np my friend - I am glad I was able to help you out. :)

OpenStudy (waheguru):

thanks

OpenStudy (asnaseer):

yw :)

OpenStudy (waheguru):

say, the root had to be by 6, then we would do 16^1/6 ?

OpenStudy (asnaseer):

correct

OpenStudy (waheguru):

Thanks

OpenStudy (asnaseer):

since:\[a^{1/b}=\sqrt[b]{a}\]

OpenStudy (asnaseer):

that is the general form

OpenStudy (waheguru):

If you have time,can I ask another similar question

OpenStudy (asnaseer):

best to post it as a new question

OpenStudy (waheguru):

sure

OpenStudy (asnaseer):

I will help if I can